OpenStudy (anonymous):

Match each percent equation with the accurate description of the equation. 1)70=0.4*b A) 70 is the part 2)a=0.3*40 B) 70 is the whole 3)40=70%*b C) 30% of the whole 4)30=p%*70 D) 40 is the part match the numbers with the matching letter.

OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!

OpenStudy (luffingsails):

Okay, here is what I would say... 1) -> A) 2) -> C) 3) -> D) 4) -> B)

OpenStudy (anonymous):

Thanks Luffingsails for your help!!!!

OpenStudy (luffingsails):

Sure, does it make sense? I am willing to help you with the wording if you want.

OpenStudy (anonymous):

thanks,

OpenStudy (anonymous):

so will you really?

OpenStudy (anonymous):

Thanks for all your help i gave you a metal! You deserve it you are very nice.

OpenStudy (luffingsails):

Sure, the question is really asking you to identify each part of a fractional equation. Fractional Part = (Fraction of the Whole) * (Whole) In that light, you can just walk through them. I can continue if you want.

OpenStudy (anonymous):

No thanks that really help a lot, Thanks for all your help i gave you a metal! You deserve it you are very nice. but can you also help me with one more?

OpenStudy (luffingsails):

Sure, ask away.

OpenStudy (anonymous):

Which of the following shows a proportion? 4/9 = 18/8 3/8=6/4 6/12=8/4 9/15=12/20

OpenStudy (luffingsails):

The key to this one is to first look at each side. Remember that a proportion like this is going to reduce to the same fraction on each side. So, we can immediately throw out the obvious ones. For instance, the first 3! The left hand sides of each of those are numbers less than 1... but the right sides are all greater than 1. So, they are not going to be equal at all. That leaves the last one. We just need to prove they are equal proportions.

OpenStudy (luffingsails):

We do that by trying to reduce or find common denominators. In this case, we can reduce: \[\frac{9}{15} = \frac{3}{5}\] \[\frac{12}{20} = \frac{3}{5}\] So, they are equal!

OpenStudy (anonymous):

Wow you are smart, let me ask another question so you can get another medal so look for it.

OpenStudy (luffingsails):

No worries about a medal. Ask here if you want.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Zach borrowed from his grandfather to buy a used car so he could travel to and from work. After working for 2 years, he has saved to pay back to his grandfather. At what interest rate did Zach borrow the money?

OpenStudy (anonymous):

Thats my question

OpenStudy (luffingsails):

hmmm... there is no information about how much he paid each month? Not sure how to approach that one.

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

he borrowd $4,000

OpenStudy (anonymous):

he saved $4,720

OpenStudy (luffingsails):

Ahh.. that is what we needed. I assume he paid his grandfather $4720? If so, then we can set up the equation. Remember the equations we worked with regarding interest above? We use those here. He paid his grandfather 4720, which is 720 dollars over the borrowed amount. 720 is also the fractional part. $4000 is the whole part. So, we set up the equation: N * Whole = Fractional Part N * 4000 = 720. Solve for N and you will have the % interest.

OpenStudy (anonymous):

thanks

OpenStudy (luffingsails):

Sure.

OpenStudy (anonymous):

so the answer is 8% right

OpenStudy (anonymous):

???????